OpenStudy (mony01):

f(x)=(x-3)^-2. Find all values of c in (1,7) such that f(7)-f(1)=f'(c)(7-1)

OpenStudy (anonymous):

Okay well we have \(f(x)=(x-3)^{-2}\) We also know that \(f'(x)=-2(x-3)^{-3}\) So we can equivocate the statement to: \[\eqalign{ &f(7)-f(1)=-2(c-3)^{-3}(6) \\ &(7-3)^{-2}-(1-3)^{-2}=-2(c-3)^{-3}(6) \\ &(4)^{-2}-(-2)^{-2}=-12(c-3)^{-3} \\ &\frac{1}{16}-\frac{1}{4}=-\frac{12}{(c-3)^3} }\] Now it is just a matter of isolating for c

OpenStudy (mony01):

@KeithAfasCalcLover how do we isolate for c?

OpenStudy (anonymous):

Ehh you would have something a little like this: \[\eqalign{ &\frac{1}{16}-\frac{1}{4}=-\frac{12}{(c-3)^2} \\ &(c-3)^2\left(\frac{1}{16}-\frac{1}{4}\right)=-12 \\ &(c-3)^2=-\frac{12}{\left(\frac{1}{16}-\frac{1}{4}\right)} \\ &c-3=\pm\sqrt{-\frac{12}{\left(\frac{1}{16}-\frac{1}{4}\right)}} \\ &c=3\pm\sqrt{-\frac{12}{\left(\frac{1}{16}-\frac{1}{4}\right)}} \\ }\]